OpenStudy (anonymous):

which ratio is equivalent to 4/16? A. 2:4 B. 6 to 20 C. 8:32 D. 12 to 64 PLEASE HELP!!

OpenStudy (kira_yamato):

Generalizing 4/16 = n/4n Plugging in n = 8, C

OpenStudy (anonymous):